# HoloLLM: Multisensory Foundation Model for Language-Grounded Human Sensing and Reasoning

### v1

@@ 2 Related Work
Extending LLMs to other modalities to form MLLMs enables them to follow human instructions based on multimodal inputs such as images [5, 8], videos [26], audio [27], and point clouds [28]. Aligning multimodal inputs with text is the key challenge for MLLMs. Existing methods primarily rely on building multimodal projectors. Typically, Linear / MLP projectors [5, 26] directly project the multimodal inputs into the LLM text space. Despite the simplicity, the number of multimodal tokens will increase significantly when high-resolution inputs are presented. To alleviate this issue, Resampler / Q-Former projectors [8, 27] utilize a fixed number of learnable queries to align the most task-relevant multimodal information with the text. Recently, hybrid projectors have been developed [6, 7], which serve as trade-offs between MLP and Q-Former projectors, to fully preserve all information within multimodal inputs while minimizing the number of tokens generated. Nevertheless, these projectors commonly need large-scale ‘modality-text’ data pairs for pre-training, which is not available for sensing modalities.
@@ 3.2 Universal Modality-Injection Projector
$\mathbf{Q}^{m}=\text{AvgPool}(Y^{m}_{CLIP})\in\mathbb{R}^{n^{\prime}_{m}\times d_{m}},$ (3)
@@
where $\mathbf{Q}^{m}_{(0)}=\mathbf{Q}^{m}$. SelfAtt($\cdot$), CrossAtt(,$\cdot$,$\cdot$,$\cdot$ , ⋅ , ⋅), and FFN($\cdot$) are self-attention, cross-attention, and feedforward layers in each block of UMIP, respectively, and $L$ is the number of blocks.
@@
Figure 3: Comparison between UMIP and other projectors: (a) Modality-Specific Projector [28, 27, 31], (b) Universal Projector [30], and (c) Universal Modality-Injection Projector (Ours).
@@
We compare UMIP with state-of-the-art multimodal projectors in Fig. 3. Specifically, most existing works [28, 27, 31] adopt modality-specific encoders and projectors (Fig. 3 (a)), which commonly requires substantial ‘modality-text’ data pairs for pre-training. As shown in Fig. 3 (b), OneLLM [30] attempts to handle various modalities via a unified framework that consists of a universal encoder and projector. However, without a dedicated design for capturing heterogeneous spatial features, the universal encoder struggles to obtain sufficiently discriminative multimodal tokens. Different from existing works, UMIP only utilizes the universal encoder to generate initial embeddings for each modality (Fig. 3 (c)). These embeddings are then progressively enhanced by fine-grained, text-aligned features from tailored encoders.
@@ V D M L W Avg V D M L W Avg
Settings Models Sources Types Action QA Action Caption
@@ V D I R W Avg V D I R W Avg
OneLLM CVPR’24 Proj 2.0 1.9 1.5 1.9 2.3 1.8 14.5 13.1 1.7 1.5 2.0 13.3
@@ 5 Conclusion and Limitations
    • C.3 Naive Multimodal Fusion for HoloLLM
@@ Appendix A Technical Details on Data Curation
In this section, we elaborate more details about the data generation process for action question answering (QA) and captions.
@@ B.3 Training Details
Settings Models Sources Types MM-Fi XRF55
@@ C.2 Complete Qualitative Results and Analysis
C.3 Naive Multimodal Fusion for HoloLLM
Modality MM-Fi Modality XRF55
@@ C.3 Naive Multimodal Fusion for HoloLLM
Table 6: Results of naive multimodal fusion on MM-Fi and XRF55 under the “Cross-Environment” setting.
We conduct a preliminary exploration of HoloLLM’s multimodal fusion capability using a naive strategy. Specifically, aligned multimodal tokens generated by UMIP are directly concatenated to achieve fusion. Multimodal reasoning is performed by prepending the fused tokens to the human instructions and passing them to the LLM. For the MM-Fi and XRF55 datasets, we consider the Vision (V), mmWave (M), WiFi (W) and Vision (V), Infrared (I), WiFi (W) modalities, respectively.
The results are summarized in Tab. 6. It shows naive multimodal fusion can enhance action recognition performance for certain modalities (highlighted in bold). However, for action QA and captioning tasks, naive multimodal fusion fails to improve performance. More advanced multimodal fusion strategies, which leverage complementary information across modalities while reducing redundancy, can be developed to enhance human perception and reasoning.

### v2

@@ 2 Related Work
Extending LLMs to other modalities to form MLLMs enables them to follow human instructions based on multimodal inputs such as images [18, 15], videos [37], audio [38], and point clouds [32]. Aligning multimodal inputs with text is the key challenge for MLLMs. Existing methods primarily rely on building multimodal projectors. Typically, Linear / MLP projectors [18, 37] directly project the multimodal inputs into the LLM text space. Despite the simplicity, the number of multimodal tokens will increase significantly when high-resolution inputs are presented. To alleviate this issue, Resampler / Q-Former projectors [15, 38] utilize a fixed number of learnable queries to align the most task-relevant multimodal information with the text. Recently, hybrid projectors have been developed [6, 17], which serve as trade-offs between MLP and Q-Former projectors, to fully preserve all information within multimodal inputs while minimizing the number of tokens generated. Nevertheless, these projectors commonly need large-scale ‘modality-text’ data pairs for pre-training, which are unavailable for sensing modalities.
@@ 3.2 Universal Modality-Injection Projector
$\mathbf{Q}^{m}=\text{AvgPool}(\mathbf{Y}^{m}_{CLIP})\in\mathbb{R}^{n^{\prime}_{m}\times d_{m}},$ (3)
@@
where $\mathbf{Q}^{m}_{(0)}=\mathbf{Q}^{m}$. SelfAtt($\cdot$), CrossAtt($\cdot,\cdot,\cdot$), and FFN($\cdot$) are self-attention, cross-attention, and feedforward layers in each block of UMIP, respectively, and $L$ is the number of blocks.
@@
Figure 3: Comparison between UMIP and other projectors: (a) Modality-Specific Projector [32, 38, 9, 42], (b) Universal Projector [11], and (c) Universal Modality-Injection Projector (Ours).
@@
We compare UMIP with state-of-the-art multimodal projectors in Fig. 3. Specifically, most existing works [32, 38, 9, 42] adopt modality-specific encoders and projectors (Fig. 3 (a)), which commonly requires substantial ‘modality-text’ data pairs for pre-training. As shown in Fig. 3 (b), OneLLM [11] attempts to handle various modalities via a unified framework that consists of a universal encoder and projector. However, without a dedicated design for capturing heterogeneous spatial features, the universal encoder struggles to obtain sufficiently discriminative multimodal tokens. Different from existing works, UMIP only utilizes the universal encoder to generate initial embeddings for each modality (Fig. 3 (c)). These embeddings are then progressively enhanced by fine-grained, text-aligned features from tailored encoders.
@@ V D I R W Avg V D I R W Avg
OneLLM CVPR’24 Proj 2.0 1.9 1.5 1.9 2.3 1.8 14.5 13.1 13.7 11.6 13.7 13.3
@@ 5 Conclusion and Limitations
    • C.3 Multimodal Fusion for HoloLLM
    • C.4 Generalization for HoloLLM
@@ Appendix A Technical Details on Data Curation
In this section, we elaborate on more details about the data generation process for action question answering (QA) and captions.
@@ B.3 Training Details
Settings Models Sources Types $M$-Fi XRF55
@@ C.2 Complete Qualitative Results and Analysis
C.3 Multimodal Fusion for HoloLLM
We conduct a preliminary exploration of HoloLLM’s multimodal fusion capability. Specifically, we first adopt a naïve fusion strategy, in which aligned multimodal tokens generated by UMIP are directly concatenated. Multimodal reasoning is achieved by prepending the fused tokens to human instructions before feeding them into the LLM. For the MM-Fi and XRF55 datasets, we consider the Vision (V), mmWave (M), WiFi (W) and Vision (V), Infrared (I), WiFi (W) modalities, respectively. The results are summarized in Tab. 6.
Modality $M$-Fi Modality XRF55
@@ C.3 Multimodal Fusion for HoloLLM
Table 6: Results of naïve multimodal fusion on MM-Fi and XRF55 under the “Cross-Environment” setting.
It shows naive multimodal fusion can enhance action recognition performance for certain modalities (highlighted in bold). However, for action QA and captioning tasks, naive multimodal fusion fails to improve performance.
Furthermore, we conduct experiments to explore more advanced adaptive fusion techniques, including weighted sum (WS) and max pooling (MP) across multimodal tokens. In the weighted-sum technique, the weights assigned to Vision and mmWave/Infrared modalities are 0.8 and 0.2, respectively.
Modality $M$-Fi Modality XRF55
HAR QA Caption HAR QA Caption
V 80.6 79.5 25.7 V 28.9 25.9 19.8
M 61.0 61.4 24.5 I 28.3 22.1 17.1
V+M (Naïve) 84.6 66.4 25.0 V+I (Naïve) 34.3 21.5 19.2
V+M (WS) 82.6 81.4 26.9 V+I (WS) 30.9 26.2 20.6
V+M (MP) 80.6 80.4 26.7 V+I (MP) 28.9 20.5 18.9
Table 7: Comparison of different multimodal fusion strategies on MM-Fi and XRF55 datasets.
As shown in Tab. 7, the weighted sum fusion strategy outperforms single-modality and naïve multimodal fusion on Action QA and captioning tasks. We believe that more advanced, learning-based multimodal fusion techniques could further improve performance. We expect HoloLLM to serve as a challenging benchmark that motivates the research community to develop deeper insights into multimodal fusion methods.
C.4 Generalization for HoloLLM
HoloLLM generalizes to new modalities more efficiently than other MLLMs due to the novel UMIP architecture. Specifically, each new modality is equipped with a lightweight, modality-specific encoder that is extensively pretrained to capture fine-grained features. The new modality is then efficiently integrated into HoloLLM by UMIP with minor data and fine-tuning.
We demonstrate that HoloLLM achieves superior data efficiency over the ‘Baseline’ model on two novel modalities: Audio [20] and UWB [21]. The results are summarized in Tab. 8.
HAR Action QA
Audio UWB Audio UWB
Baseline 37.71 9.46 30.73 11.33
HoloLLM 63.94 16.77 54.17 16.45
Table 8: Results of HAR and Action QA tasks on novel Audio and UWB modalities.
Compared to the ‘Baseline’ model, HoloLLM can be efficiently generalized to novel modalities with minor data and fine-tuning.